he following information relates to a capital project that will last for five years (dollars are in thousands). The discount rate is 10%. (Round all numbers to whole numbers.) Purchase of land and building ($2,450) Training of staff ($50) Annual cash inflow from new facility $2,100 Annual cash outflow from new facility ($1,800) Salvage value of facility in five years $500. what is the initial outflow? What is the annual net inflow - please show work
Answer #1

1. Initial Cash outlay = Purchase cost + Training of staff

=2450+50

=2500

2. Annual net inflow

Years 1 to 4 =Annual cash inflow - Annual cash outflow

=2100-1800

=300

Year 5 =( Annual cash inflow - Annual cash outflow ) + Salvage value

=300+500

=800